OpenStudy (anonymous):

For which value of x does the graph of y = 2x^2 − 7x + 6 cross the x-axis?

OpenStudy (anonymous):

The answer comes out to be \[-(-7) \pm \sqrt{1} \div{4}\]

OpenStudy (lgbasallote):

you used quadratic formula?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Igb, please explain how to do this.

OpenStudy (lgbasallote):

your quadratic formula is right...you will end up with \[\large x = \frac{7 \pm 1}{4}\] agree?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

therefore \[\large x = \frac{7+1}{4}\] and \[x = \frac{7-1}{4}\] therefore... \[x = \frac{8}{4}\] and \[x = \frac{6}{4}\] simplifying further will get you \[x = 2\] and \[x = \frac{3}{2}\] so those are the values of x where the graph crosses x-axis

OpenStudy (anonymous):

No no..

OpenStudy (lgbasallote):

no?

OpenStudy (anonymous):

\[-\frac {3}{2}\]

OpenStudy (anonymous):

You see where you messed up?

OpenStudy (lgbasallote):

where?

OpenStudy (anonymous):

I don't know. but on my options to choose from \[\frac {3}{2} \]

OpenStudy (anonymous):

Isn't on there.

OpenStudy (anonymous):

So you messed up somewheres.

OpenStudy (lgbasallote):

if you do this by factoring... \[2x^2 - 7x + 6 = (2x - 3)(x-2)\] if i equate each factor to 0.. \[2x - 3 = 0\] and \[x - 2 = 0\] therefore... \[2x = 3\] and \[x = 2\] if i simplify 2x = 3 i will get \[x = \frac{3}{2}\]

OpenStudy (lgbasallote):

factoring results in same values

OpenStudy (anonymous):

well my only options to choose from are 2. 3. \[-\frac {2}{}\]

OpenStudy (anonymous):

Oh my god.

OpenStudy (anonymous):

UGHGHGH

OpenStudy (anonymous):

2. 3. -2 / 3. and -3 / 2.

OpenStudy (lgbasallote):

then choose 2...x = 2 also isn't it

OpenStudy (anonymous):

So 2?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

Okay I'll tell you if you're right.

OpenStudy (anonymous):

Hmm. I made 7 out of 10, but the 3 I missed, i did on my own. You got it right.

OpenStudy (lgbasallote):

yay :)

OpenStudy (anonymous):

Are you good with biology?

OpenStudy (lgbasallote):

nope..but i believe @Diyadiya is good in bio...i believe her course is bio math

OpenStudy (anonymous):

Mie is too.

OpenStudy (anonymous):

Well, I have a few more math questions if you're willing to help.

OpenStudy (lgbasallote):

sure..just post them on the left side and i'll be happy to pop in if i know how it is

OpenStudy (anonymous):

ok